OpenStudy (anonymous):

For this intgeral could i let u=x^3+1, then let x=3root (u-1)

hero (hero):

What integral?

OpenStudy (anonymous):

\[\int\limits_{}^{} (x^5)(\sqrt[4]{x^3+1})\]

hero (hero):

wait a sec

hero (hero):

Satellite, if you want to take a stab at it you can

OpenStudy (anonymous):

then \[u-1=x^3\] so \[x^5=(u-1)^{\frac{5}{3}}\] does that help?

OpenStudy (anonymous):

oh no i see split \[x^5=x^2\times x^3\] maybe that will work

hero (hero):

I attempted that approach

hero (hero):

the splitting approach

OpenStudy (anonymous):

before i cheat let me try that \[\int \sqrt[4]{x^3+1}x^2x^3dx\] \[

OpenStudy (anonymous):

\[u=x^3+1\] \[du=3x^2\] \[x^3=u-1\] should work right?

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

get \[\frac{1}{3}\int \sqrt[4]{u}(u-1)du\]

hero (hero):

Looks good

OpenStudy (anonymous):

oh sorry i wrote before i saw your post.

hero (hero):

I was on the right track. No I hadn't posted anything yet

OpenStudy (anonymous):

it cool let me see if i can replicate

OpenStudy (anonymous):

okay i got: (4/27)(x^3+1)^9/4 -(4/5)(x^3+1)^5/4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

please check!!